# Alchemist: Towards the Design of Efficient Online Continual Learning System

### v1

@@ Abstract.
Continual learning has emerged as a promising solution to refine models incrementally by leveraging user feedback, thereby enhancing model performance in applications like code completion, personal assistants, and chat interfaces. In particular, online continual learning — iteratively training the model with small batches of user feedback — has demonstrated notable performance improvements. However, the existing practice of segregating training and serving processes forces the online trainer to recompute the intermediate results already done during serving. Such redundant computations can account for 30%–42% of total training time.
In this paper, we propose Alchemist, to the best of our knowledge, the first online continual learning system that efficiently reuses intermediate results computed during serving to reduce redundant computation with minimal impact on the serving latency or capacity. Alchemist introduces two key techniques: (1) minimal activations recording and saving during serving, where activations are recorded and saved only during the prefill phase to minimize overhead; and (2) offloading of serving activations, which dynamically manages GPU memory by freeing activations in the forward order, while reloading them in the backward order during the backward pass. Evaluations with the ShareGPT dataset show that compared with a separate training cluster, Alchemist significantly increases training throughput by up to 1.72x, reduces up to 47% memory usage during training, and supports up to 2x more training tokens — all while maintaining negligible impact on serving latency.
@@ 1. Introduction
Large Language Model (LLM) services, widely adopted in modern cloud ecosystems, power applications like chatbots, search engines, recommendation systems, and API offerings. The market is projected to grow from \$6.5B (2024) to \$140.8B (2033), as 92% of Fortune 500 companies integrate cloud-hosted LLMs for generative AI workflows (TechTarget Editorial, 2023; Topics, 2024; Company, 2023). While training LLM is highly resource and computation intensive when feeding a large amount of data, serving the trained models is often even more expensive due to requirements on low latency, which causes less efficient resource utilization. To keep the model performant and maintain a profitable business, it is often required to periodically retrain the model. While most resources are allocated to serving for the most of the time, such periodically training could also require a significant amount of resources.
Among the training methods, continual learning (Gabriel, 2020; Zhang et al., 2023a; Sun et al., 2020; Zhang et al., 2023b; Cheng et al., 2024; Gabriel, 2020; Zan et al., 2022; Song et al., 2023; Cossu et al., 2022; Gururangan et al., 2020) has become increasingly popular because it allows models to be refined frequently without the need to retrain them from scratch. It incrementally adapts based on up-to-date real world information and interactions with users or models that mimic user preferences (Guo et al., 2024; Lee et al., 2024b; DeepSeek-AI et al., 2025; Xiang et al., 2025). For example, chatbots with web search capability can learn from the up-to-date searched context (e.g., news, financial data, et al.) to provide the latest information in future responses. Or, code completion applications allow users to accept or reject suggested code, and this user feedback can be used to allow models to continuously learn their preference, providing better and more personalized code completion.
Current common practice in a continual learning typically involves collecting large batches of data (e.g., user prompts, user feedback, or model generated labels) on a single fixed model over an extensive period of time (e.g., several months). However, many prior works have shown that online continual learning2.4), which iteratively trains the model based on a small batch of user feedback or model-generated labels collected each time the model updated, can greatly improve model performance (Guo et al., 2024; Qi et al., 2024a; Gao et al., 2024b; Shaikh et al., 2024; Xu et al., 2024; Carta et al., 2024; Gaven et al., 2024). Unlike traditional continual learning, online continual learning can gather more accurate user feedback on the models that has been retrained for some iterations, leading to superior performance benefits.
@@
In this context, we observe a significant gap between the modern need for LLM services to update iteratively and frequently and the conventional design practices of LLM systems. As illustrated in Figure 1, existing systems typically separate training from serving to maintain quality and low latency. This separation occurs both temporally (i.e., data collected during serving is only used for training after a considerable delay) and spatially (i.e., training is carried out on machines that are distinct from those used for serving). However, such hard separation loses substantial opportunities to improve the model quality and reduce computational costs.
@@
We observe that, with such design, online continual learning trainer needs to repeat the same computation that already had conducted by the serving process, and this redundant computation often contributes to a significant portion of the entire computation for training.
In reality, depending on the training methods, we measure the trainer would spend 30%–42% of the total time for the computations that already been done during serving. This observation reveals a key insight used in our paper: the intermediate results produced during serving can and should be reused in the forward pass in training procedure.
@@
If the training and serving processes are separate on different machines as existing systems designed, the intermediate results from serving need to be either recomputed during training or sent to training machines via network. The latter potentially leads to prohibitive overhead in network transmission due to the size of the intermediate results, undermining the goal of reusing serving activations.
@@
An intuitive and straightforward design is to remove the hard separation between serving and training and co-locate serving and training on the same machine, leveraging GPU multiplexing techniques (e.g., either temporal sharing or spatial sharing). So, as once serving produces activations, they can be used directly and promptly by the trainer, without any transmission delay to send activations across machines.
@@
However, activation reusing can still be non-trivial even if we remove the transmission overhead by co-locating training with serving. First of all, although reusing activations can reduce training latency, to perform backward pass with the reused activations, the serving procedure needs to record and store the activations produced by each layer and the forward computation graph (§4.1.1). This procedure incurs extra latency overhead to the forward passes during serving, which could potentially violate serving latency SLO. Concretely, such latency overhead can be up to 35% when storing and recording activations and computation graphs. Secondly, in a normal serving procedure, each layer’s intermediate results may be overwritten by the next layers’ intermediate results to reduce overall memory footprint. Hence, storing the each layer’s intermediate results for reusing in training also incurs extra memory overhead and severely limits the serving capacity due to the large size of activations.
@@
To address the above challenges, we propose Alchemist, to the best of our knowledge, the first efficient system that reuses serving activations in training to increase online training throughput, with minimal impact on both serving latency and capacity. Alchemist primarily entails two techniques:
@@
    Minimal activation recording during serving: To minimize the activations recording overhead during LLM serving, we only record the activations when processing user input processing, i.e., prefill phase, and disable it when generating new tokens, i.e., decoding phase. This can be effective because in real-world scenarios (anon8231489123, 2025), the total decoding time is often the majority of the end-to-end time. Furthermore, in the case where activations are required for both input and model response, Alchemist reuses the KV cache (one kind of intermediate results) generated by prefill in serving to avoid activations recording for the entire generations procedures while still managed to reuse part of the activations calculated during serving (§4.1.1).
@@
    Offloading of serving activations: To guarantee serving capacity, when the serving requires more GPU memory based on an offline memory capacity profiler, Alchemist frees the serving activations in a as needed fashion. Based on the insight that the activations of earlier layers are needed at last during backward pass, Alchemist frees activations in forward order and loads them in the reverse order of the layers, maximizing the overlap between backward computation and activations loading.
We compare Alchemist with the baseline that separates the training and serving processes on different machines, and show that on a popular dataset, ShareGPT (anon8231489123, 2025):
@@
    Alchemist increases the training throughput by 1.26x–1.72x, across two different example continual learning methods.
@@ 2.2. Parameter Efficient Fine-Tuning (PEFT)
With the rapid growth in LLMs’ size, the cost of continual adapting models with the new information also skyrockets due to the requirements of large number of GPUs. PEFT methods (Li and Liang, 2021; Lester et al., 2021; Kopiczko et al., 2024; Qiu et al., 2023; Liu et al., 2023a; Blau et al., 2024; Liu et al., 2023b) mitigate training costs by freezing the base model and only updating low-rank adapters hence reducing the memory and computation requirements. Consequently, these methods, especially LoRA (Hu et al., 2021), have become the most favorable methods for fine-tuning LLMs.
@@ 2.3. Continuous Adaptation
Continual Pretraining:  Models ingest fresh data (news, user queries, code commits) to update factual knowledge (Sun et al., 2020; Zhang et al., 2023b; Cheng et al., 2024; Gabriel, 2020; Zan et al., 2022; Cossu et al., 2022; Gururangan et al., 2020). For example, a financial LLM might retrain daily on earnings reports to improve stock analysis.
@@ 2.4. Online Continuous Learning
Typical procedure in continuous learning often aggregates data from serving side over weeks or months. But these delayed updates face challenges, for instance: concept drift occurs as user preferences shift, such as changes in coding style trends; factual obsolescence arises when news, APIs, or regulations become outdated.
Ideally, we want the model to be continuously trained on almost a real-time stream of information and feedback. This means that, instead of infrequent, large updates, model is frequently retrained and updated with small batches of latest. Users (or LLMs that mimic users’ preferences), in turn, provide the latest feedback or information based on the response generated by the new version to further update and improve the model’s performance iteratively.
This cyclic and iterative training process is often referred as online learning or online continuous learning (Guo et al., 2024; Qi et al., 2024a; Gao et al., 2024b; Shaikh et al., 2024; Xu et al., 2024; Carta et al., 2024; Gaven et al., 2024). Unlike the traditional offline continuous learning procedure which update the model based on one large batch of data at one time, online continuous learning pipeline emphasizes its nature in iterative improvement over time based on a micro-batch of newest information and feedback collected each time after the model is updated.
@@ 2.5.1. Reusing Activations
Furthermore, in an online continuous learning scenario, since user prompts or user feedback will be the training input (or part of the input), the forward passes during serving have already calculated the activations we need for backpropagation. Such forward passes can cost as much as 43% of the total training time, shown in red color in Fig 3. Hence, intuitively, it is more efficient to reuse the activations calculated from serving forward pass during backpropagation rather than re-calculate it again.
@@
However, this schema can be challenging and suboptimal in the design of separated training and serving cluster. As the model size (i.e., number of parameters) rapidly increases, so does the size of activations as well the number of activations. Attempt to reusing the activations calculated by serving cluster on a separate training cluster requires transmitting a large amount of data (i.e., activations) over network, and could incur significant transmission overhead. It is likely that such overhead can eventually surpass the benefits of reusing the activations (i.e., transmission latency is higher than the recomputation latency). Consequently, this design then has to face a dilemma where it has to choose either transmit activations with high network cost or recompute the activations with high computation cost, both of which are inefficient.
Obviously, a more resource-efficient strategy here is to co-locate training and serving on the same infrastructure to avoid either transmission or re-computation cost. This co-location design enables a powerful key insight: model activations computed during serving can be re-used to reduce or even remove the forward passes in the training loop.
@@ 3. Design Requirements & Assumption
Reusing activations for online continual learning and leveraging idle computation cycles is an appealing strategy for cost efficiency. However, designing such a system is non-trivial. We envision that the system should, at a minimum, meet the following design requirements:
@@
    A2: GPUs’ computation resources are overprovisioned to ensure service quality in peak or unexpected spikes in workload.
@@ 4. Alchemist Design
4.1. Activation Cache
To fulfill R1 of reusing serving activations for training, the first step is to record the required activations and cache them for later training use. Although this may sound straightforward, under the requirements of R2 and R3, this is rather non-trivial.
To simplify the design process, we will address the requirements individually by temporarily setting aside the capacity requirement from R3.
4.1.1. Recording Activations
After the prefill phase, one query may require several hundreds or thousands decodes to generate the full response and return to users. If we record and save all the activations and the backpropagation computation graph, it could incur serious latency overhead from the autograd frameworks (e.g., torch.enable_grad), shown in Fig 5. This clearly violates the latency requirement from R2.
However, depending on the type of continuous training and how their loss functions are orchestrated, not all activations calculated during serving may or may not be needed during training.
Continual pre-train:  For continual pretrain, we only need to store the activations from the prefill phase during serving. When pretraining a LLM, the loss function is designed to measure the averaged difference between the probability distribution of the next generated token and the actual next token averaged by attending on all the tokens from the corpus.
Continual pretrain, in terms of the loss functions, generally follows the same method as pretraining. In this case, the prompt or the searched context are the corpus we hope the model learns from, loss calculation and backpropagation only requires the activations calculated when attending the input (i.e., prefill phase). This tells that we only need to enable activation recording during prefill phase. Since prefill only happens once and is the minority of the total serving time, such overhead then can be amortized in the end-to-end latency as decode proceeding.
@@ 4.1.1. Recording Activations
Figure 5. Latency overhead of activations recording. Due to the cost in recording and saving activations and computation graphs, autograd frameworks like torch.autograd could bring up 21% overhead to the prefill phase and 35% overhead to each forward pass in the decode phase (i.e., 35% increment in each token’s generation time). If enabled for each token generated, it will significantly prolong the serving latency, violating our latency requirement. §4.1.1
Continual preference alignment:  In the case of continual preference alignment, this problem can be complicated. In preference alignment algorithms like DPO and its other variations, since we directly compare the chosen and rejected responses, we typically need to make two forward passes during loss calculation, one with the chosen response and one with the rejected response (both prefixed with prompt).

### v2

@@ Abstract.
Continual learning has become a promising solution to refine large language models incrementally by leveraging user feedback. In particular, online continual learning — iteratively training the model with small batches of user feedback — has demonstrated notable performance improvements. However, the existing practice of separating training and serving processes forces the online trainer to recompute the intermediate results already done during serving. Such redundant computations can account for 30%–42% of total training time.
In this paper, we propose Alchemist, to the best of our knowledge, the first online continual learning system that efficiently reuses serving activations to increase training throughput. Alchemist introduces two key techniques: (1) recording and storing activations and KV cache only during the prefill phase to minimize latency and memory overhead; and (2) smart activation offloading and hedging. Evaluations with inputs of varied token length sampled from ShareGPT dataset show that compared with a separate training cluster, Alchemist significantly increases training throughput by up to 1.72x, reduces up to 47% memory usage during training, and supports up to 2x more training tokens — all while maintaining negligible impact on serving latency.
@@ 1. Introduction
Large Language Model (LLM) services, widely adopted in modern cloud ecosystems, power applications, such as chatbots, search engines, recommendation systems, and API offerings. The market is projected to grow from \$6.5B (2024) to \$140.8B (2033), as 92% of Fortune 500 companies integrate cloud-hosted LLMs for generative AI workflows (TechTarget Editorial, 2023; Topics, 2024; Company, 2023). To ensure these models remain accurate and continue to enhance user experiences, organizations must regularly retrain them to learn the latest data and refine performance.
One of the most popular and promising training approaches is continual learning (Gabriel, 2020; Zhang et al., 2023a; Sun et al., 2020; Zhang et al., 2023b; Cheng et al., 2024; Gabriel, 2020; Zan et al., 2022; Song et al., 2023; Cossu et al., 2022; Gururangan et al., 2020), due to its capability of allowing models to be refined frequently, without training from scratch. It incrementally adapts the models with up-to-date real-world information and interactions with users or models that mimic user preferences (Guo et al., 2024; Lee et al., 2024b; DeepSeek-AI et al., 2025; Xiang et al., 2025). For example, chatbots with web searching capability can learn from the newest searched context (e.g., latest news, financial data, et al.) to provide the latest information in future responses. Or, code completion applications allow users to accept or reject suggested code, and this user feedback can be used to allow models to continuously learn their preference, providing better and more personalized code completion and debug suggestions.
One of the most popular continual learning approach, online continual learning2.4), iteratively trains the model based on a small batch of user feedback or model-generated labels collected each time the model updated (Guo et al., 2024; Qi et al., 2024a; Gao et al., 2024b; Shaikh et al., 2024; Xu et al., 2024; Carta et al., 2024; Gaven et al., 2024).
@@
In this context, we observe a significant gap between the modern need for iterative and frequent updates in LLM services and the conventional paradigm for LLM training. As illustrated in Figure 1, existing systems typically separate training from serving to maintain low latency for serving. This separation occurs both temporally (i.e., data collected during serving is only used for training after a considerable delay) and spatially (i.e., training is carried out on machines that are distinct from those used for serving). However, such hard separation loses substantial opportunities to reduce computational costs.
@@
An important and interesting observation we made is that, with such design, online continual learning trainer needs to repeat the same computation that has already been conducted by the serving process, and this redundant computation often contributes to a significant portion of the entire computation for training. In reality, the trainer would spend 30%–42% of the total time for the computations that have already been done during serving, depending on the loss functions.
This observation reveals a key insight of our approach: the intermediate results (i.e., activations) generated during serving can be directly re-used in the forward pass of training. Specifically, we reuse these activations from serving to eliminate redundant forward-pass computations.
@@
However, if the training and serving processes are separate on different machines as existing systems designed, to reuse the serving actions, they need to be sent to the training machines via the network, posing large network transmission overhead given the large size of activations.
@@
An intuitive and straightforward design is to remove the hard separation between serving and training and co-locate serving and training on the same machine, leveraging GPU multiplexing techniques (e.g., either temporal sharing or spatial sharing). So, once there is idle time in the serving process (e.g., during off-peak hours at night time), the serving activations can be used directly by the trainer, without any transmission delay across machines.
A potential concern around this design is whether the serving workload offers sufficient idle resources. Nowadays, the serving resource is typically over-provisioned  (Kubernetes, 2025) to ensure end users’ experience (e.g.latency service-level objectives (SLOs) guarantee). This indicates that a significant amount of idle resources is always available and can be harvested for training usage. Another concern is whether the newly collected serving data is safe to be immediately incorporated into model updates and whether the updated model can be immediately deployed to users. To address potential safety risks, multiple studies (Kim and Lee, 2024; Inan et al., 2023; Morimura et al., 2024; Ousidhoum et al., 2021; Ji et al., 2024) have proposed safeguards to ensure only safe data can be trained by the model. Many other works (Guo et al., 2024; Qi et al., 2024a; Gao et al., 2024b; Shaikh et al., 2024; Xu et al., 2024; Carta et al., 2024; Gaven et al., 2024; Wang et al., 2023b) focus on ensuring that the updated models neither drift too far away from the pre-trained model, losing its generality, nor generate unethical or unlawful responses.
@@
However, there is still a system-level challenge that remains unsolved. Specifically, co-locating serving and training on the same machine may incur extra overhead in saving and reusing the serving activations. First of all, although reusing activations can reduce training latency, the serving process needs to record and store the activations produced by each layer and the forward computation graph, to be used in the backward pass for gradient computation (§4.1). This procedure incurs additional latency overhead to the forward passes during serving, which could potentially violate serving latency SLO. Concretely, such latency overhead can be up to 35% when storing and recording activations and computation graphs. Secondly, in a normal serving procedure, each layer’s activations are overwritten by the next layers’ activations to reduce overall memory footprint (Kwon et al., 2023). Hence, storing each layer’s activations for reusing in training also incurs extra memory overhead and severely limits the serving capacity due to the large size of activations.
@@
To address the above challenges, we propose Alchemist, to the best of our knowledge, the first efficient system that reuses serving activations in training to increase online training throughput, with minimal impact on both serving latency and capacity. Alchemist entails two techniques:
@@
    Minimal activation recording during serving: To minimize the activations recording overhead during LLM serving, we only record the activations when processing user input, i.e., prefill phase, and disable it when generating new tokens, i.e., decoding phase. This can be effective because in real-world scenarios (anon8231489123, 2025), the total decoding time is often the majority of the end-to-end time. Furthermore, in the case where activations are required for both input and model response, Alchemist reuses the KV cache (one kind of intermediate results) generated by prefill in serving to avoid activations recording for the entire generation procedures while still managing to reuse part of the activations calculated during serving (§4.1).
@@
    Offloading of serving activations: To guarantee serving capacity, when the serving requires more GPU memory based on an offline memory capacity profiler, Alchemist frees the serving activations in a as needed fashion. Based on the insight that the activations of earlier layers are needed at last during the backward pass, Alchemist frees activations in forward order and loads them in the reverse order of the layers, maximizing the overlap between the backward computation and activations loading.
We compare Alchemist with the baseline that separates the training and serving processes on different machines, and show that on input with varied token length sampled from a popular dataset, ShareGPT (anon8231489123, 2025):
@@
    Alchemist increases the training throughput by 1.26x–1.72x, across two different example continual learning methods with three types of LLMs.
@@ 2.2. Parameter Efficient Fine-Tuning (PEFT)
With the rapid growth in LLMs’ size, the cost of continually adapting models with the new information also skyrockets due to the requirements of a large number of GPUs. PEFT methods (Li and Liang, 2021; Lester et al., 2021; Kopiczko et al., 2024; Qiu et al., 2023; Liu et al., 2023a; Blau et al., 2024; Liu et al., 2023b) mitigate training costs by freezing the base model and only updating low-rank adapters with much smaller parameter sizes, hence reducing the memory and computation requirements. Consequently, these methods, especially LoRA (Hu et al., 2021), have become the most favorable methods for fine-tuning LLMs.
@@ 2.3. Continuous Adaptation
Continual Pretraining:  Models ingest fresh data (news, user prompts, code commits) to update factual knowledge (Sun et al., 2020; Zhang et al., 2023b; Cheng et al., 2024; Gabriel, 2020; Zan et al., 2022; Cossu et al., 2022; Gururangan et al., 2020). For example, a financial LLM might retrain daily on earnings reports to improve stock analysis.
@@ 2.4. Online Continuous Learning
Typical procedure in continuous learning often aggregates data from the serving side over weeks or months, known as offline learning. But these delayed updates face challenges, for instance: concept drift occurs as user preferences shift, such as changes in coding style trends; factual obsolescence arises when news, APIs, or regulations become outdated.
Ideally, we want the model to be continuously trained on almost a real-time stream of information and feedback. This means that, instead of infrequent, large updates, model is frequently retrained and updated with small batches of latest. Users (or LLMs that mimic users’ preferences), in turn, provide the latest feedback or information based on the response generated by the updated version to further update and improve the model’s performance iteratively.
This cyclic and iterative training process is often referred to as online continuous learning (Guo et al., 2024; Qi et al., 2024a; Gao et al., 2024b; Shaikh et al., 2024; Xu et al., 2024; Carta et al., 2024; Gaven et al., 2024). Unlike the traditional offline continuous learning procedure which trains the models based on a single large batch of stationary data for every iteration, online continuous learning pipeline emphasizes its nature in iterative improvement over time based on a micro-batch of newest information and feedback collected each time after the model is updated.
@@ 2.5.1. Reusing Activations
Furthermore, in an online continuous learning scenario, since user prompts or user feedback will be the training input (or part of the input), the forward passes during serving have already calculated the activations we need for backpropagation. Such forward passes can cost as much as 43% of the total training time, shown in red color in Fig 3. Hence, intuitively, it is more efficient to reuse the activations calculated from the serving forward pass during backpropagation rather than re-calculate them again.
@@
However, this schema can be challenging and suboptimal in the design of separated training and serving clusters. As the model size (i.e., number of parameters) rapidly increases, so does the size of activations as well as the number of activations. Attempt to reuse the activations calculated by the serving cluster on a separate training cluster requires transmitting a large amount of data (i.e., activations) over the network and could incur significant transmission overhead. It is likely that such overhead can eventually surpass the benefits of reusing the activations (i.e., transmission latency is higher than the recomputation latency). Consequently, this design then has to face a dilemma where it has to choose either to transmit activations with high network cost or to recompute the activations with high computation cost. Both of which are inefficient.
Obviously, a more resource-efficient strategy here is to co-locate training and serving on the same infrastructure to avoid either transmission or re-computation costs. This co-location design enables a powerful key insight: model activations computed during serving can be re-used to reduce or even remove the forward passes in the training loop.
@@ 3. Design Requirements & Assumption
Reusing activations for online continual learning and leveraging idle computation cycles is an appealing strategy for cost efficiency. However, designing such a system is non-trivial. Due the intensive computation and large memory requirements from training, the key challenge lies in how to preserve serving latency and capacity while performing training at the same time. Hence, we envision that the system should, at a minimum, meet the following design requirements:
@@
    A2: GPUs’ computation and memory resources are overprovisioned to ensure service quality in peak or unexpected spikes in workload.
@@ 4. Alchemist Design
Figure 5. Latency overhead of activations recording. Due to the cost in recording and saving activations and computation graphs, autograd frameworks like torch.autograd could bring up 21% overhead to the prefill phase and 35% overhead to each forward pass in the decode phase (i.e., 35% increment in each token’s generation time). If enabled for each token generated, it will significantly prolong the serving latency, violating our latency requirement. §4.1
4.1. Record and cache activation
To simplify the design process, we set aside R3 for now by assuming we have more than enough GPU memory to co-locate serving and training. To fulfill R1 of reusing serving activations for training, the first step is to record the required activations and cache them for later training use. Although this may sound straightforward, under the requirements of R2, this can be rather non-trivial. Due to the overhead from the autograd frameworks (e.g., torch.enable_grad), shown in Figure 5, thoughtlessly enabling activation recording during the generation process can severely impact serving latency, clearly violating R2.
Continual pre-train:  In the case of continual pre-train, model is learning from users’ input prompts or the searched context. Hence, loss calculation and backpropagation only require the activations calculated when processing the input (i.e., prefill phase). This tells that we only need to enable activation recording during prefill phase. Since prefill only happens once and is the minority of the total serving time, such overhead can then be amortized in the end-to-end latency as decode proceeding.
Continual preference alignment:  In the case of continual preference alignment, this problem can be complicated. In preference alignment algorithms like DPO and its other variations, since they directly compare the chosen and rejected responses, it would require the full response from the model to calculate the loss. Hence, naïvely reusing all the activations of the model generated responses (either it is the chosen or the rejected one) would require us to record and save the activations until the end of generation during serving. This could introduce prohibitive overhead.